The average weight of a medium size egg is found to be 3.2 oz, with a standard deviation of 0.13 oz. What is the heaviest egg (to 2 decimal places) that we are 95% confident is still medium size?
3.2 oz
3.33 oz
3.45 oz
3.46 oz

Answer by stanbon(75887)

The average weight of a medium size egg is found to be 3.2 oz, with a standard deviation of 0.13 oz. What is the heaviest egg (to 2 decimal places) that we are 95% confident is still medium size?
----
Draw a normal curve.
Sketch a 95% area centered at 3.2 oz
---
Find the z-value of the upper limit of that area.
invNorm(0.975) = 1.96
-----
Find the corresponding weight value::
x = z*s + u
x = 1.96*0.13+3.2 = 3.45 oz
---------------
Cheers,
Stan H.
-------------


3.2 oz
3.33 oz
3.45 oz
3.46 oz